Find f 0 (0) for the function f(x) = sin−1 (x^2 + x) + 5x.
Answer by Alan3354(69443)

Find f 0 (0) for the function f(x) = sin-1(x^2 + x) + 5x.
---------------
Sub 0 for x:
f(0) = sin-1(0+0) + 5*0
= sin-1(0)